OpenStudy (anonymous):

f(x) = 4x + 1 and g(x) = 3x + 2, what is f(g(x))

OpenStudy (anonymous):

this means exactly what it looks like. Plug g(x) into f(x) It looks like f(g(x)) f(3x +2) f(3x+2)= 4(3x+2) + 1 f(3x +2)= 12x + 8 +1 f(g(x)) = 12x + 9 I hope that helps!

OpenStudy (anonymous):

Just remember that g(x) is just some number! So plugging g(x) into f(x) is just like plugging any other number into f(x)